OpenStudy (anonymous):

Find all solutions to the equation. sin^2x + sin x = 0

OpenStudy (anonymous):

a^2+a=a(a+1)

OpenStudy (helder_edwin):

exactly \[ \large 0=\sin^2x+\sin x=\sin x(\sin x+1) \]

OpenStudy (anonymous):

so Sinx=0, pi, 2pi and -1?

OpenStudy (anonymous):

sin(x)=0 gives you x=kpi where k=0, 1, 2, 3, ... sin(x)+1=0 gives you x=k3pi/2 where k=0, 1, 2, 3, ...

OpenStudy (anonymous):

ah okay i got it now thank you (: